The price of Profile, Inc., stock will be either $82 or $104 at the end of the year. Call options are available with one year to expiration. T-bills currently yield 3 percent.

  

a.

Suppose the current price of the stock is $93. What is the value of the call option if the exercise price is $78 per share? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

b. Suppose the current price of the stock is $93. What is the value of the call option if the exercise price is $88 per share? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

a

Upmove (U)= High price/current price=104/93=1.1183
Down move (D)= Low price/current price=82/93=0.8817
Risk neutral probability for up move
q = (e^(risk free rate*time)-D)/(U-D)
=(e^(0.03*1)-0.8817)/(1.1183-0.8817)=0.62874
Call option payoff at high price (payoff H)
=Max(High price-strike price,0)
=Max(104-78,0)
=Max(26,0)
=26
Call option payoff at low price (Payoff L)
=Max(Low price-strike price,0)
=Max(82-78,0)
=Max(4,0)
=4
Price of call option = e^(-r*t)*(q*Payoff H+(1-q)*Payoff L)
=e^(-0.03*1)*(0.62874*26+(1-0.62874)*4)
=17.31

b

Upmove (U)= High price/current price=104/93=1.1183
Down move (D)= Low price/current price=82/93=0.8817
Risk neutral probability for up move
q = (e^(risk free rate*time)-D)/(U-D)
=(e^(0.03*1)-0.8817)/(1.1183-0.8817)=0.62874
Call option payoff at high price (payoff H)
=Max(High price-strike price,0)
=Max(104-88,0)
=Max(16,0)
=16
Call option payoff at low price (Payoff L)
=Max(Low price-strike price,0)
=Max(82-88,0)
=Max(-6,0)
=0
Price of call option = e^(-r*t)*(q*Payoff H+(1-q)*Payoff L)
=e^(-0.03*1)*(0.62874*16+(1-0.62874)*0)
=9.76